(10) 2. Solve the homogeneous equation by making the substitution y = xv y x + 2y 2x + y > 0.
Answer #1

do polve homageneous equation by making substite ton Y=vn. y = anty x+2y 7 2+24 dy da anty Given that xv equation with repect2 dv 6) V dv t 1-v2 da 71. 1-V? T I IT Now we 1-V dv ln / It will solve all in leqal.fon 1. His put 1-4²=t -2y = dt -dt v dv.well 13 v.] *** toote (S1-v)(3-1) un una & C. Jan Titu In la ta 312 all team are in Un} than the Songlant In Info + Gov)317 fDo the complele Solution 13. f =2 C n² 1- h V